

********************************************************************
UARS WEEKLY STATUS REPORT - 31 MAR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
We are working on a version of the production software which more
gracefully handles incomplete scans, several of which were gathered
during the recent period of unpredicted telemetry outages.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. AXIS-2
was deactivated on March 5 and AXIS-1 deactivation occurred on March 13. This
week on March 27, AXIS-1 was again activated. At the present time, PEM energy
deposition is not being generated. Recent data configuration changes have
affected PEM software causing multiple failures. These failures are being
reviewed by the PEM team. In order to correct for new telemetry requirements,
PEM will need to update processing code with a patch delivery. Updated HEPS
calibration parameters were finalized and are ready for delivery. Updated MEPS
quality calibration files are being generated which are used during Level 3
product generation. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE was powered on today, March 31, 2000. All engineering parameters are
within nominal operating region.
***********
MLS
***********
MLS was taken out of 'standby' mode on 26 March 2000 and operated until 30
March 2000. The objective of this operation was to obtain information on
Arctic vortex ozone chemistry. MLS was operated in a mode where limb scans
were performed at ~50N-80N latitudes.
The MLS electronics on (those needed to support the 205 GHz radiometer and
measurements of ClO, HNO3, O3, UTH, CH3CN) operated perfectly. No slips
of the antenna scan mechanism occurred this observing period.
MLS is now back in its Standby Mode.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG distributed the Instrument Operations Schedule for the next
yaw period: 25 April to 4 June 2000.
Routine processing of instrument Daily Activity Plans continues with
no problems.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HAL (4.82) PROD (1997) (1R-1C=0L)
1 day SEN (5.0) PROD (1992) (1R-1C=0L)
The CDHF will load February 2000 current day DAAC data in April 2000.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM,
Solstice, and WINDII.
- WINDII V5.11A tested successfully on the Rehost System. Thanks to WINDII
for validating the output product in a timely manner.
- Testing HRDI V4.4 on the Rehost System continues.
- Continue to validate Level 0 Ingest and NRT.
- Contacting suppliers of correlative data to establish interfaces to the
rehost system.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 March, 2000 (Orbit 46548, GMT Day 078) through 24 March
2000 (Orbit 46653, GMT Day 084).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
---------- ---- ----- --------
HALOE 079/2342 46563 Off
The beta angle ranged between the angles of 11.3 degrees to -0.884 degrees for
this report period. The beta angle is now increasing toward a maximum angle
+49.0 degrees on 08 April 2000 (DOY 099).
Spacecraft battery 1, 2, and 3 performance data is not available for this
report period.
The clock error ranged from -10.0 to 11.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
Forward-to-Reverse Yaw Around maneuver (number 87) was successfully executed
in Orbit 46603 beginning 081/16:00:59 and completing at 081/16:38:09 on 21
March 2000. The SSPP Instruments (ACRIM, SOLSTICE, SUSIM) were enabled as well
as HRDI and PEM in day only operations. Due to the SA position at 269 degrees,
the effective night length of the spacecraft after the maneuver was
approximately 50 minutes. Post maneuver analysis of data indicate that during
the Yaw Around, the end-of-night load bus voltage was 24.12 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were approximately 22.5
and 23.0 percent respectively. The next (reverse-to-forward) YAW is scheduled
for 25 April 2000 (DOY 116).
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
FOT = 06
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 00/021 - 2287.496849
Transponder B: 00/060 - 2287.499207
GROUND SYSTEM OPERATION
Two (2) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/083 show the observatory data loss to be 566 hours, 38
minutes, 01.921 seconds (an increase of 16 minutes 30.984 seconds since the
last report period).
The increase of 16 minutes 30.984 seconds occurred over three (3) days:
06 minutes 03.976 seconds on Day 077. Reason unknown.
10 minutes 17.344 seconds on Day 078. Reason unknown.
00 minutes 09.664 seconds on Day 079. Reason unknown.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT provided coordinates of the High Gain Antenna (HGA) to Flight Dynamics
Facility. They have some new software that they are running and they needed an
accurate measurement of the HGA's position with respect to the center of mass.
The FOT was unable to locate accurate coordinates but were able to determine a
coarse estimate. FDF appeared to be satisfied with what was provided. The FOT
will continue to look for accurate UARS dimensions for future reference.
The FOT continued work on the safehold procedures for UARS. This week RTS-52
was coded into the TMON reaction listing which was the last step of the
development of the safehold algorithms. The remaining steps are all related to
the configuration of the simulator and testing
MEETINGS
March 23, 2000 (DOY083): The FOT met with the Flight Software team. Two
individuals have been assigned the development of three TDRSS operations. They
have previous experience from the GRO mission which was modified to three TDRSS
operations. The FOT is planning to meat with them next week and provide any
documentation that they request.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for April 25, 2000 (DOY 116). This
maneuver will be a Reverse to Forward Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 March 2000
ORBIT AOS NO SUBSYS REASON
----- --- --- ------ -------
46551 078/0441 00-111 NBTR NBTR OPS
46563 079/2342 00-112 HALOE POWER OFF
46568 079/0739 00-113 NBTR NBTR OPS
46602 081/1509 00-114 OBC YAW MANEUVER
46623 083/0106 00-115 OBC Clk Adj fr 0.00 to -20.6 msec/day
46639 084/0150 00-116 OBC Clk Adj fr 20.6 to 0.00 msec/day
46641 084/0508 00-117 NBTR NBTR OPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 March 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON R-LO PTS PULSE BUS CURRENT MON 46638-40
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None to report this period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 March 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46623 083/01:06 (23 Mar 00) from 0.00 to -20.6 msec/day
46639 084/01:56 (24 Mar 00) from -20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 March 2000
17 Mar 2000 - GMT Day 077 Beta = 11.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- --------- ---------- ---- ------- ------
1 -44.8/-67.2 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +22.4/-5.6 +5.79/+5.15 +30.0/-12.8 24.2 1.030 22.0
3 +50.4/+11.2 +4.51/+3.87 +28.8/-12.8 24.1 1.027 21.7
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ----- ------- ------
Battery performance data not available for this report period.
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_____________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 March 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-005 FOT 46600 081/1149 14170 GENERIC LATE ACQ
00-005 FOT 46614 082/1021 14170 GENERIC LATE ACQ
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No closure information received
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-005 FOT see list March 00 14170 Generic Late Acq.
SEVNTFW00/12